Circular No. 7689

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001dt IN UGC 12558 W. D. Li, University of California at Berkeley, reports the discovery

by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 18.2) in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on Aug. 23.4 and 24.4 UT. The new object is located at α = 23^{h}22^{m}19^{s}.17, δ = +50^{o}09′25′′.3 (equinox 2000.0), which is 1′′.3 west and 8′′.0 south of the nucleus of UGC 12558. A KAIT image taken on Aug. 14.4 UT showed nothing at this position (limiting mag ∼ 19.5).

COMETS C/2001 P1 AND C/2001 P2 (SOHO) Further to IAUC 7667, astrometry (measured by D. Hammer, reduced

by B. G. Marsden) and orbital elements are given on the MPECs cited below for two additional sungrazing comets found by T. Scarmato and S. Hönig, respectively, on SOHO website images:

Comet  2001 UT  α_{2000}  δ_{2000}  MPEC
C/2001 P1 Aug. 4.971  8^{h}48.^{m}5  +16^{o}12′  2001-P22
C/2001 P2  15.221  9 22.7  +13 02  2001-Q02

COMET 39P/2001 P3 (OTERMA) Y. R. Fernandez, University of Hawaii, reports his recovery of comet

39P on CCD frames obtained on Aug. 13 (see below) with the 2.2-m University of Hawaii reflector, confirmatory images being obtained on Aug. 20 and 21 by K. J. Meech and J. Pittichová. The object, a point source, was located ∼ 2′ from the prediction by B. G. Marsden on MPC 34423 (ephemeris on MPC 42373). Meech then succeeded in locating the comet on her CCD frames from 1999 May 9 and July 15. M. A. Kadooka and J. M. Bauer assisted, and the measurements by Meech are given on MPEC 2001-Q35. The recovery also confirms the correctness of positions tentatively measured by G. V. Williams from images obtained by D. C. Jewitt, J. X. Luu, and C. A. Trujillo on 1998 May 1 and 22. MPEC 2001-Q35 also includes orbital elements from 227 observations (1942–2001) and a revised ephemeris. Last observed in Aug. 1962, comet 39P passed 0.095 AU from Jupiter on 1963 Apr. 12, after which q increased from 3.4 to 5.5 AU and P from 7.9 to 19 years (with T = 1983 June 18 and 2002 Dec. 22).

2001 UT  α_{2000}  δ_{2000}  m_{2}
Aug. 13.42834  23^{h}29^{m}40.^{s}16  −2^{o}37′11′′.5  22.0
2001 August 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green